# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
  • A second property of the uniform distribution is that it can be sampled by a sequentially-invertible algorithm, in fact, by a $1$-invertible circuit sequence. To see this, observe that to sample uniformly at random from $\{0,1,\cdots,2^{\beta+1}-1\}$, all we need to do is to break the sampled number into its summation as powers of $2$, and in the $i$-th circuit for $i\in[\beta]$, we add (or not) the number $2^{i}$. Summing random powers of $2$ up to power $\beta$ exactly gives a uniformly random distribution in $\{0,1,\cdots,2^{\beta+1}-1\}$. At the end we can deterministically move the distribution to center around zero, by subtracting $2^{\beta}$.
  • To see why the sequence is sequentially invertible, observe that for every $i\in[\beta]$, for every possible randomness for the circuit $C_{i}$ (which is just one bit $b_{i}\in\{0,1\}$), the fixed-randomness circuit $C_{i}\left(\cdot;b_{i}\right)$ always adds modulo $2^{\beta+1}$ the same number $b_{i}\cdot 2^{i}$. By recalling Fact 2.1 once again,
To conclude, in our reduction, the circuit $C^{0}$ samples a uniformly random lattice vector $\mathbf{v}$, and then adds an approximate discrete Gaussian vector $\mathbf{e}$, where the approximation is done by summing i.i.d. samples from the uniform distribution. The circuit $C^{1}$ is the same, only that we add $\mathbf{t}$ in the end. Both circuits are shown to be $1$-sequentially invertible. We refer the reader to Section 7 for the full details of the proof, including the analysis of the total variation distance between the circuits.
@@ 3 Preliminaries
In this section we introduce previously existing notions and definitions from the literature that are relevant to this work.
We use the following notions and notations, most of which are standard in quantum computation and complexity.
@@
  • In this work, when we use the notation $\log(\cdot)$ we implicitly refer to the base-$2$ logarithm function $\log_{2}(\cdot)$.
  • In this work, the only non-standard notation will be for iterated matrix multiplication. Specifically, our default convention for iterated matrix multiplication, using the notation $\prod$, will be by enumerating from right to left (unlike the usual, left to right). Specifically, for a natural number $m\in\mathbb{N}$ and square complex matrices $\{M_{i}\}_{i\in[m]}$ ($\forall i\in[m]:M_{i}\in\mathbb{C}^{n\times n}$ for some $n\in\mathbb{N}$), we define:
@@
  • We follow the standard Dirac notations for quantum states in quantum information processing. For a variable $x$ (i.e., notation) that denotes a classical binary string, i.e., $x\in\{0,1\}^{*}$, the notation $|{x}\rangle$ refers to the pure quantum state of $|x|$ qubits, which equals that standard basis element of the space $\mathbb{C}^{2^{|x|}}$ that corresponds to the classical string $x$. That is,
@@
  • For two classical random variables $D_{0}$, $D_{1}$ over binary strings of finite length $k$, the total variation distance (sometimes referred to as statistical distance) between $D_{0}$ and $D_{1}$ is defined as follows:
@@
  • For two classical random variables $D_{0}$, $D_{1}$ over binary strings of finite length $k$, the fidelity between $D_{0}$ and $D_{1}$ is defined as follows:
@@
  • For every pair of classical random variables $D_{0}$, $D_{1}$ over binary strings of finite length $k$, the following relation between the fidelity and total variation distance between distributions is known:
@@
  • Unless specifically noted otherwise, for any variable $u$ (not necessarily denoting a classical string), the notation $|{u}\rangle$ denotes a unit vector in complex vector space, of some finite dimension. That is, the notation usually refers to a quantum state and is thus a normalized vector, unless explicitly noted otherwise. Also, for any $n\in\mathbb{N}$, an $n$-qubit state is always a unit vector in $\mathbb{C}^{2^{n}}$.
@@
  • A PPT algorithm is a classical probabilistic polynomial-time Turing machine. It is a known fact in computational complexity theory that a PPT is equivalent to a uniform family of polynomial-sized classical probabilistic circuits. We will use these two notions interchangeably throughout this work as our model for efficient classical computation.
@@
  • For $n\in\mathbb{N}$, an $n$-qubit general circuit (or $n$-qubit quantum algorithm) is a quantum circuit $C$ that can perform the constant-qubit quantum gates $\{X,Z,P,H,T,CNOT,SWAP,Tof\}$, execute measurement gates and add ancillary qubits registers containing $|{0^{k}}\rangle$.
@@
  • For $n\in\mathbb{N}$ and an $n$-qubit unitary $U$, a unitary circuit for $U$, denoted $C_{U}$ (or $U$, when it is clear from the context whether we are referring to the matrix or circuit version of $U$) is an $n$-qubit, $k$-ancilla (for some $k\in\left(\mathbb{N}\cup\{0\}\right)$) quantum circuit that implements the unitary $U$.
  • Representing permutations: For $m\in\mathbb{N}$, our way for presenting a permutation $\sigma\in S_{m}$ in the form of a binary string in this paper, is by using $m\cdot\lceil\log_{2}(m)\rceil$ bits, interpreted as a list of $m$ numbers in $[m]$ – an ordering of the set $[m]$ according to the permutation $\sigma$. Specifically, the number in the $i$-th packet of $\lceil\log_{2}(m)\rceil$ bits, is $\sigma^{-1}(i)$.
@@ 3.1 Standard Notions from Computational Complexity Theory
[…]
@@
[…]
@@
[…]
The complexity class BQP represents all computational problems that have a feasible solution on a quantum computer, and more broadly (and in short), BQP represents humanly-feasible quantum computation. The writing BQP stands for Bounded-Error Quantum Polynomial-Time, and defined as follows (which is one out of many equivalent definitions of this class).
@@
[…]
@@
[…]
@@
As part of this work we show how a strengthening of the model of quantum computation solves the graph isomorphism problem. To this end, we define (simple and undirected) graphs and then the graph isomorphism problem. First, a (simple and undirected) graph with $n$ vertices is simply a set of $n$ points, implicitly numbered $[n]$, where some pairs of points have edges between them. The edges are are lines, or connections, between a pair of points $i,j\in[n]$, $i\neq j$. In this work when we refer to graphs, they are implicitly simple and undirected.
@@
[…]
@@
In case the pair are isomorphic, then for a permutation $\sigma\in S_{n}$ on the set $[n]$, denote by $\sigma\left(G\right)$ the graph generated by rearranging the vertices of $G$ with accordance to the permutation $\sigma$. then, the following distributions are identical:
@@
We turn to the definition of the gap-closest vector problem ($\mathsf{GapCVP}$). We define lattices and then the $\mathsf{GapCVP}$ problem.
For a basis $\mathbf{B}$ of $\mathbb{R}^{n}$, a lattice with basis $\mathbf{B}$ is the infinite set of all integer linear combinations of the basis vectors in $\mathbf{B}$. In this work, in order to not deal with real-values vectors and finite precision, we consider only lattice bases with integer coordinates, i.e., bases of $\mathbb{R}^{n}$ from $\mathbb{Z}^{n}$. The computational hardness of these problems are polynomially equivalent.
@@
[…]
@@ 3.2 Statistical and Algorithmic Tools
Throughout this work we use the Swap Test [${}^{+}$97, BCWDW01] algorithm, which, among other things, can be used to distinguish between quantum states as a function of their inner product as vectors.
@@
The central limit theorem (CLT), one of the fundamental theorems in statistics, tells us that when we take $N\in\mathbb{N}$ independent and continuous random variables $X_{N},\cdots,X_{N}$, the sum of the samples approaches the Gaussian distribution, as $N$ goes to infinity. The CLT does not tell us the rate at which the samples approach the Gaussian distribution, as this depends, among other things, on the probability density functions of the summed variables $\{X_{i}\}_{i\in[N]}$. The Berry-Esseen Theorem is a quantified version of the central limit theorem – it bounds the Wasserstein-1 distance between the sum $S:=\sum_{i\in[N]}X_{i}$ and the Gaussian distribution, as a function of $N$ and the density functions of the variables $\{X_{i}\}_{i\in[N]}$. The original formulation of the Berry-Essen Theorem is insufficient for our needs, because (1) In this paper we are working with discrete random variables and not continuous, and (2) We want a bound on the total variation distance which implies physical statistical indistinguishability, and for this, the Wasserstein-1 distance is not enough. We will use the following strengthening Theorem from [CGS10]:
@@
[…]
@@ 4.1 Quantum Computation with Computable Order Interference
Let $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary circuits which all operate on the same number of qubits $n\in\mathbb{N}$ and let $|{\psi}\rangle$ an $n$-qubit quantum state.
@@
  • $R$ is an $n$-qubit quantum register,
  • and $\lambda\in\mathbb{N}$ is a natural number.
@@
$\sum_{x\in\{0,1\}^{n^{\prime}}}\alpha_{x}\cdot|{\psi_{x}}\rangle_{R}|{x}\rangle_{R^{\prime}}\enspace.$
@@
$\min_{x\in\{0,1\}^{n^{\prime}}}\left(\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\right)\enspace.$
@@
$\sum_{x\in\{0,1\}^{k^{\prime}}}\alpha_{x}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle\right)_{R}|{x}\rangle_{R^{\prime}}\enspace,$
@@
Claim 4.1 (Corollary - Uniform OI on pure states).
@@
$\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\enspace,$
and with (at least) the same probability, the state in $R$ after the oracle call is the normalization of the uniform OI state, $\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$.
@@
Let $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary circuits which all operate on the same number of qubits $n\in\mathbb{N}$, let $|{\psi}\rangle$ an $n$-qubit quantum state, and let $U_{O}$ a $k$-qubit unitary circuit on registers $\left(O,A\right)$ where $O$ has $m\cdot\lceil\log\left(m\right)\rceil$ qubits and $A$ has the rest.
@@
$\sum_{x\in\{0,1\}^{n^{\prime}}}\alpha_{x}\cdot|{\psi_{x}}\rangle_{R}|{x}\rangle_{R^{\prime}}\enspace.$
@@
$\min_{x\in\{0,1\}^{n^{\prime}}}\left(\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}+\frac{1}{\lambda}}\right)\enspace.$
@@
$\sum_{x\in\{0,1\}^{n^{\prime}}}\alpha_{x}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{x}}\rangle,U_{O}\right)_{R}|{x}\rangle_{R^{\prime}}\enspace,$
@@
Then, the success probability of the oracle call $\mathcal{O}_{\textbf{OI}}\left(\lambda,\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$ is $\geq 1-\frac{1}{k}$.
@@
[…]

### v2

@@ 2.4 Reducing $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
  • A second property of the uniform distribution is that it can be sampled by a sequentially-invertible algorithm, in fact, by a $1$-invertible circuit sequence. To see this, observe that to sample uniformly at random from $\{0,1,\cdots,2^{\beta+1}-1\}$, all we need to do is to break the sampled number into its summation as powers of $2$, and in the $i$-th circuit for $i\in[\beta]$, we add (or not) the number $2^{i}$. Summing random powers of $2$ up to power $\beta$ exactly gives a uniformly random sample in $\{0,1,\cdots,2^{\beta+1}-1\}$. At the end we can deterministically move the distribution to center around zero, by subtracting $2^{\beta}$.
  • To see why the circuit sequence is invertible, observe that for every $i\in[\beta]$, for every possible randomness for the circuit $C_{i}$ (which is just one bit $b_{i}\in\{0,1\}$), the fixed-randomness circuit $C_{i}\left(\cdot;b_{i}\right)$ always adds modulo $2^{\beta+1}$ the same number $b_{i}\cdot 2^{i}$. By going back to Fact 2.1, it follows that the fixed-randomness $C_{i}\left(\cdot;b_{i}\right)$ is efficiently computable and invertible, as needed.
To conclude, the circuit sequence $C^{0}$ samples a uniformly random lattice vector $\mathbf{v}$, and then adds an approximate discrete Gaussian vector $\mathbf{e}$, where the approximation is done by summing i.i.d. samples from the uniform distribution. The circuit sequence $C^{1}$ is the same, only that we add $\mathbf{t}$ in the end. Both circuit sequences are shown to be $1$-sequentially invertible. We refer the reader to Section 7 for the full details of the proof, including the analysis of the total variation distance between the output distributions of the circuits.
@@ 3 Preliminaries
This section contains previously existing notions and definitions from the literature that are relevant to this work, mostly from quantum computation and computational complexity theory.
@@
  • When we use the notation $\log(\cdot)$ we implicitly refer to the base-$2$ logarithm function $\log_{2}(\cdot)$.
  • The only non-standard notation will be for iterated matrix multiplication. Specifically, our default convention for iterated matrix multiplication, using the notation $\prod$, will be by enumerating from right to left (unlike the usual, left to right). Specifically, for a natural number $m\in\mathbb{N}$ and square complex matrices $\{M_{i}\}_{i\in[m]}$ ($\forall i\in[m]:M_{i}\in\mathbb{C}^{n\times n}$ for some $n\in\mathbb{N}$), we define:
@@
  • We follow the standard Dirac notations for quantum states in quantum information processing. For a variable $x$ (i.e., notation) that denotes a classical binary string, i.e., $x\in\{0,1\}^{*}$, the notation $|{x}\rangle$ refers to the pure quantum state of $|x|$ qubits, which equals the $x$-th standard basis element of the space $\mathbb{C}^{2^{|x|}}$. That is,
@@
  • For two classical random variables (i.e., distributions) $D_{0}$, $D_{1}$ over binary strings of finite length $\leq k$, the total variation distance (sometimes referred to as statistical distance) between $D_{0}$ and $D_{1}$ is defined as follows:
@@
  • For two classical random variables (i.e., distributions) $D_{0}$, $D_{1}$ over binary strings of finite length $\leq k$, the fidelity between $D_{0}$ and $D_{1}$ is defined as follows:
@@
  • For every pair of classical random variables $D_{0}$, $D_{1}$ over binary strings of finite length $\leq k$, the following relation between the fidelity and total variation distance between distributions is known:
@@
  • Unless specifically noted otherwise, for any variable $u$ (not necessarily denoting a binary string), the notation $|{u}\rangle$ denotes a unit vector in complex vector space, of some finite dimension. That is, the notation usually refers to a quantum state and is thus a normalized vector, unless explicitly noted otherwise. Also, for any $n\in\mathbb{N}$, an $n$-qubit state is always a unit vector in $\mathbb{C}^{2^{n}}$.
@@
  • A PPT algorithm is a classical probabilistic polynomial-time Turing machine. It is a known fact in computational complexity theory that a PPT is equivalent to a uniform family of polynomial-sized classical probabilistic circuits. We will use these two notions interchangeably throughout this work as our model for efficient probabilistic classical computation.
@@
  • For $n\in\mathbb{N}$, an $n$-qubit general circuit (or $n$-qubit quantum algorithm) is a quantum circuit $C$ that can perform the constant-sized quantum gates $\{X,Z,P,H,T,CNOT,SWAP,Tof\}$, execute measurement gates and add ancillary qubits registers containing $|{0^{k}}\rangle$.
@@
  • For $n\in\mathbb{N}$ and an $n$-qubit unitary transformation $U$, a unitary circuit for $U$, denoted $C_{U}$ (or denoted $U$ by overloading notation, when it is clear from the context whether we are referring to the matrix or circuit version of $U$) is an $n$-qubit, $k$-ancilla (for some $k\in\left(\mathbb{N}\cup\{0\}\right)$) unitary quantum circuit that implements the unitary $U$.
  • Representing permutations: For $m\in\mathbb{N}$, our way for presenting a permutation $\sigma\in S_{m}$ in the form of a binary string in this paper, is by using $m\cdot\lceil\log(m)\rceil$ bits, interpreted as a list of $m$ numbers in $[m]$ – an ordering of the set $[m]$ according to the permutation $\sigma$. Specifically, the number in the $i$-th packet of $\lceil\log_{2}(m)\rceil$ bits, is $\sigma^{-1}(i)$.
@@ 3.1 Standard Notions from Computational Complexity Theory
Let $\{0,1\}^{*}:=\bigcup_{i\in\left(\mathbb{N}\cup\{0\}\right)}\{0,1\}^{i}$ the set of all binary strings, of all lengths. A promise problem $\prod:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ is given by two (possibly infinite) sets $\prod_{\text{YES}}\subseteq\{0,1\}^{*}$, $\prod_{\text{NO}}\subseteq\{0,1\}^{*}$ such that $\prod_{\text{YES}}\cap\prod_{\text{NO}}=\emptyset$.
@@
Let $\prod:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ a promise problem. We say that $\prod$ is a decision problem if $\prod_{\text{YES}}\cup\prod_{\text{NO}}=\{0,1\}^{*}$.
@@
Let $\prod:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ a decision problem. We say that $L$ is the language induced by the decision problem $\prod$ if $\prod_{\text{YES}}=L$.
The complexity class BQP represents all computational problems that have a feasible solution on a quantum computer, and more broadly (and informally), BQP represents humanly-feasible quantum computation. The writing BQP stands for Bounded-Error Quantum Polynomial-Time, and defined as follows (which is one out of many equivalent definitions of this class).
@@
The complexity class BQP is the set of promise problems $\prod:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ that are solvable by a quantum polynomial-time algorithm. Formally, such that there exists a QPT $M$ such that for every $x\in\prod_{\text{YES}}$, $M(x)=1$ with probability at least $p(x)$, and if $x\in\prod_{\text{NO}}$, $M(x)=0$ with probability at least $p(x)$, such that
$\forall x\in\left(\prod_{\text{YES}}\cup\prod_{\text{NO}}\right):p(x)\geq 1-2^{-\mathsf{poly}\left(|x|\right)}\enspace,$
@@
  • $\prod_{\text{YES}}$ is the set of pairs of circuits $\left(C_{0},C_{1}\right)$ such that $\left\lVert D_{0}-D_{1}\right\rVert_{TV}\leq a(n)$,
  • $\prod_{\text{NO}}$ is the set of pairs of circuits $\left(C_{0},C_{1}\right)$ such that $\left\lVert D_{0}-D_{1}\right\rVert_{TV}>b(n)$,
@@
As part of this work we show how a strengthening of the model of quantum computation solves the graph isomorphism problem. To this end, we define (simple and undirected) graphs and then the graph isomorphism problem. First, a (simple and undirected) graph with $n$ vertices is a set of $n$ points, implicitly numbered $[n]$, where some pairs of points have edges between them. The edges are directionless lines, or connections, between a pair of points $i,j\in[n]$, $i\neq j$. In this work when we refer to graphs, they are implicitly simple and undirected.
@@
  • $\prod_{\text{YES}}$ is the set of pairs of graphs $\left(G_{0},G_{1}\right)$ such that $G_{0}\simeq G_{1}$.
  • $\prod_{\text{NO}}$ is $\{0,1\}^{*}\setminus\prod_{\text{YES}}$.
@@
In case the pair is isomorphic, then for a permutation $\sigma\in S_{n}$ on the set $[n]$, denote by $\sigma\left(G\right)$ the graph generated by rearranging the vertices of $G$ with accordance to the permutation $\sigma$. Then, the following distributions are identical:
@@
We turn to the definition of the gap closest vector problem ($\mathsf{GapCVP}$). We define lattices and then the $\mathsf{GapCVP}$ problem. For a basis $\mathbf{B}$ of $\mathbb{R}^{n}$, a lattice with basis $\mathbf{B}$ is the infinite set of all integer linear combinations of the basis vectors in $\mathbf{B}$. In this work, in order to not deal with real-valued vectors and finite precision, we consider only lattice bases with integer coordinates, i.e., bases of $\mathbb{R}^{n}$ from $\mathbb{Z}^{n}$. The computational hardness of these problems are polynomially equivalent.
@@
  • $\prod_{\text{YES}}$ is the set of triplets $\left(\mathbf{B},\mathbf{t},d\right)$ such that $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)\leq d$.
  • $\prod_{\text{NO}}$ is the set of triplets $\left(\mathbf{B},\mathbf{t},d\right)$ such that $\Delta\left(\mathcal{L}_{\mathbf{B}},\mathbf{t}\right)>d\cdot g(n)$.
@@ 3.2 Statistical and Algorithmic Tools
We use the Swap Test [BBD+97, BCWDW01] algorithm, which, among other things, can be used to distinguish between quantum states as a function of their inner product as vectors.
@@
The central limit theorem (CLT), one of the fundamental theorems in statistics, tells us that when we take $N\in\mathbb{N}$ independent and continuous random variables $X_{1},\cdots,X_{N}$, the uniform average of the samples approaches the Gaussian distribution, as $N$ goes to infinity. The CLT does not tell us the rate at which the samples approach the Gaussian distribution, as this depends, among other things, on the probability density functions of the summed variables $\{X_{i}\}_{i\in[N]}$. The Berry-Esseen Theorem is a quantified version of the central limit theorem – it bounds the Wasserstein-1 distance between the sum $S:=\sum_{i\in[N]}X_{i}$ and the Gaussian distribution, as a function of $N$ and the density functions of the variables $\{X_{i}\}_{i\in[N]}$. The original formulation of the Berry-Essen Theorem is insufficient for our needs, because (1) In this paper we are working with discrete random variables and not continuous, and (2) We want a bound on the total variation distance which implies physical statistical indistinguishability, and for this, the Wasserstein-1 distance is not enough. We will use the following strengthening Theorem from [CGS10]:
@@
  • For every $i\in[N]$, the variance of $X_{i}$, $\sigma_{i}^{2}:=\text{Var}\left(X_{i}\right):=\mathbb{E}\left(X_{i}^{2}\right)-\mathbb{E}\left(X_{i}\right)^{2}$.
  • The variance of the sum $S$, $\sigma^{2}:=\text{Var}\left(S\right)$, which equals $\sum_{i\in[N]}\sigma_{i}^{2}$, due to the random variables being independent.
@@ 4.1 Quantum Computation with Computable Order Interference
Let $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary transformations which all operate on the same number of qubits $n\in\mathbb{N}$ and let $|{\psi}\rangle$ an $n$-qubit quantum state.
@@
  • $R$ is an $n$-qubit quantum register, and
  • $\lambda\in\mathbb{N}$ is a natural number.
@@
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot|{\psi_{y}}\rangle_{R}|{y}\rangle_{R^{\prime}}\enspace.$
@@
$\min_{y\in\{0,1\}^{n^{\prime}}}\left(\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{y,x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{y,x,\sigma}|}\right)\cdot\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\right)\enspace,$
where for $y\in\{0,1\}^{n^{\prime}}$, $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi_{y}}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{y,x,\sigma}|{x}\rangle\enspace.$
@@
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle\right)_{R}|{y}\rangle_{R^{\prime}}\enspace,$
@@
- The unitaries $\{U_{i}\}_{i\in[m]}$ can also be given in the form of controlled oracle access. In that case, the complexity $|U_{i}|$ for $i\in[m]$ is the complexity of making an oracle call to the $(n+1)$-qubit unitary which is the controlled $U_{i}$.
@@
Corollary 4.1 (Corollary - Uniform OI on pure states).
@@
$\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{x,\sigma}|}\right)\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\enspace,$
where for $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,\sigma}|{x}\rangle\enspace.$
and with (at least) the same probability, the state in $R$ after the oracle call is the normalization of the uniform OI state, $\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$. In case the call fails, the state in $R$ may be arbitrary.
@@
Let $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary transformations which all operate on the same number of qubits $n\in\mathbb{N}$, let $|{\psi}\rangle$ an $n$-qubit quantum state, and let $U_{O}$ a $k$-qubit unitary circuit on registers $\left(O,A\right)$ where $O$ has $m\cdot\lceil\log\left(m\right)\rceil$ qubits and $A$ has the rest.
@@
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot|{\psi_{y}}\rangle_{R}|{y}\rangle_{R^{\prime}}\enspace.$
@@
$\min_{y\in\{0,1\}^{n^{\prime}}}\left(\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{y,x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{y,x,\sigma}|}\right)\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}+\frac{1}{\lambda}}\right)\enspace,$
where for $y\in\{0,1\}^{n^{\prime}}$, $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi_{y}}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{y,x,\sigma}|{x}\rangle\enspace.$
@@
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)_{R}|{y}\rangle_{R^{\prime}}\enspace,$
@@
- The unitaries $\{U_{i}\}_{i\in[m]}$ can also be given in the form of controlled oracle access. In that case, the complexity $|U_{i}|$ for $i\in[m]$ is the complexity of making an oracle call to the $(n+1)$-qubit unitary which is the controlled $U_{i}$.
@@
Then,
$\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\geq 1-\frac{1}{k}\enspace.$
@@
The complexity class $\textbf{BQP}^{\textbf{OI}}$ is the set of promise problems $\prod:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ such that there exists $M^{\textbf{OI}}$ a QPT with access to an OI oracle $\mathcal{O}_{\textbf{OI}}$ (as in Definition 4.4), such that for every $x\in\prod_{\text{YES}}$, $M^{\textbf{OI}}(x)=1$ with probability at least $p(x)$, and if $x\in\prod_{\text{NO}}$, $M^{\textbf{OI}}(x)=0$ with probability at least $p(x)$, such that
$\forall x\in\left(\prod_{\text{YES}}\cup\prod_{\text{NO}}\right):p(x)\geq 1-2^{-\mathsf{poly}\left(|x|\right)}\enspace,$